ſi, if 0 St<T, y + 4y = 10, if a St< 0. y(0) = 0, y(0) = 0. 9
Answer #1

given y + 4y= { if of te T if Tato y(o)=0, y (o)=0 By definition of Laplace transform Lq fLt)} = d eft f(t) Lt Formula:- {y1

-PT (p+A) J (P) - - 2 EP P ý (P) = ЕРТ elporta) P(P²+4) Now Applying Inverse Laplace transform PT P p²+4 .) P [ { } (0)} = -22

3

If you satisfy with my answer please give?